OpenStudy (anonymous):

Evaluate. (16 + 32 ÷ 4)(21 – 6 • 3)

OpenStudy (noseboy908):

The most important thing here is PEMDAS. Are you familiar with this method, or any other, of remembering the order of operations?

OpenStudy (anonymous):

the answer is 72

OpenStudy (anonymous):

Parenthases Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

32/4=8+16=24 . 6.3=18, 21-18=3, 24.3=72..

OpenStudy (noseboy908):

Correct! So, noting that the first thing we have to take care of is the parenthesis, let's do that. In the first set given, what's the first thing we have to do?

OpenStudy (anonymous):

add

OpenStudy (noseboy908):

No, look again at your list.

OpenStudy (anonymous):

nope. gotta divide

OpenStudy (noseboy908):

Jay, please stop.

OpenStudy (anonymous):

stop what? helping?

OpenStudy (anonymous):

i was about to say Division i thought i was wrong

OpenStudy (anonymous):

tatty you are doing great. would you prefer me to help you?

OpenStudy (noseboy908):

You're right, don't worry. Now, having divided 32/4, now we add it to 16. That equals what/

OpenStudy (noseboy908):

*?

OpenStudy (anonymous):

i just need help i dont care who its from and 24

OpenStudy (anonymous):

okay I totslly got you, but I will let him take over, I don't wanna be rude. good luck girly!

OpenStudy (noseboy908):

Correct Tatyy, so now we'll keep that number in mind for later, and move to the second set of parenthesis. According to PEMDAS, which is the first part we do in it?

OpenStudy (anonymous):

multiplication?

OpenStudy (noseboy908):

Correct. So, please do that. What does that part equal to?

OpenStudy (anonymous):

im multiplying 24 and 6....im sorry im home schooled they never really explained

OpenStudy (noseboy908):

Why are you multiplying 24 and 6?

OpenStudy (anonymous):

ooo no 3 i think

OpenStudy (anonymous):

nevermind i get it

OpenStudy (noseboy908):

Ok, so what does the second set of parenthesis simplify to?

OpenStudy (anonymous):

18..?

OpenStudy (noseboy908):

That's what the second two integers combine to equal, but if you understand, what does the whole second set of parenthesis equal to?

OpenStudy (anonymous):

21-18= 3 then 24 x 3 = 72..... i hope i did this right

OpenStudy (noseboy908):

Don't worry, you're completely correct! That's your final answer.

OpenStudy (anonymous):

I said that 10 minutes ago, but I suppose I should have explained it..

OpenStudy (anonymous):

you just said the answer

OpenStudy (anonymous):

i know. i said i should have explained it.

OpenStudy (anonymous):

oh ..thank you

OpenStudy (anonymous):

you are welcome.